A small sample of coal, when completely converted into energy, releases 3.75 * 10 ^ 34j What is the initial mass of coal, assuming the final mass is 0? 


1
Expert's answer
2021-01-13T11:38:11-0500

 Coal has a specific heating value


"2.9\times 10^7J/kg"

If the final mass is taken as 0, then we get


"\frac{3.75\times 10^{34}}{2.9\times 10^7}=1.29\times 10^{28} kg"